Cost of goods manufactured

Primare Corporation has provided the following data concerning last month's manufacturing operations.

Purchases of raw materials $30,000
Indirect materials included in manufacturing overhead $5,000
Direct labor $58,000
Manufacturing overhead applied to work in process $87,000
Underapplied overhead $4,000

Inventories Beginning Ending
Raw materials $12,000 $18,000
Work in process $56,000 $65,000
Finished goods $35,000 $42,000

Requirement 1:
Calculate the cost of goods manufactured for the month.
Requirement 2:
Calculate the cost of goods sold for the month.
Answer #1
DM used = DM Beg + DM Purch – DM End
= 12000 + 30000 – 18000 = 24000

COGM = 24000+56000+58000+87000-65000 = 160,000

COGS = 35000+160000-42000 = 9000
Answer #2

Solution:

1) Schedule of Cost of Goods manufactured

Primare Corporation

Schedule of Cost of Goods Manufactured

Direct Materials:

Beginning Raw Material

12,000

(+) Raw Material purchases during the month

30,000

Total Raw Material available

42,000

(-) Ending Raw material

(18,000)

Raw Material used in production

24,000

(-) Indirect Material included in manufacturing Overhead (Since indirect material is a part of manufacturing overhead and in this question it is already included in Manufacturing Overhead, hence deduced from here)

(5,000)

19,000

(+) direct labor

58,000

(+) Manufacturing overhead applied to work in process

87,000

Total Manufacturing Costs

164,000

(+) Beginning Work In Process

56,000

220,000

(-) Ending Work in Process

(65,000)

Cost of Goods manufactured

155,000

2) Schedule of Cost of Goods Sold

Primare Corporation

Schedule of Cost of Goods Sold

Finished Goods Inventory, beginning

35,000

(+) Cost of Goods Manufactured

155,000

Goods available for sale

190,000

(-) Finished Goods Inventory, Ending

(42,000)

Unadjusted cost of goods sold

148,000

(+) Under-applied Overhead (Under-applied overhead means the overhead was less applied i.e. Actual Overheads are higher than Applied Overhead, hence added to the cost of goods sold)

4,000

Adjusted Cost of Goods Sold

152,000

Hope the above calculations, working and explanations are clear to you and help you in understanding the concept of question.... please rate my answer...in case any doubt, post a comment and I will try to resolve the